OpenStudy (jane11509):

the population of jamesburg is 12,000 Junction City's is 3/25 of Jamesburg's What is junction city's population? A. 1440 B. 3992 C. 4008 D. 100,000

OpenStudy (adajiamcneal):

question

OpenStudy (itsbribro):

what is the question?

OpenStudy (jane11509):

sorry i couldn't copy paste it for some reason

OpenStudy (adajiamcneal):

Population of Junction city = 3/25 times the population of Jamesburg Population of Junction city = 3/25 *12,000 =3*480 =1440

OpenStudy (jane11509):

Thank you ^-^

OpenStudy (adajiamcneal):

welcome :)